

********************************************************************
UARS WEEKLY STATUS REPORT - 09 FEB 2001
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, taking solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. We continue to gather enough data to meet all the SOLSTICE
science objectives during this period of reduced tape recorder operations.
We began day only operations on February 9th and will continue this mode
through February 10th.
This week, SOLSTICE personnel met with the FOT and off-line engineers at
GSFC regarding the December operations anomaly. We discussed exactly how
SOLSTICE is operated with its microprocessor loads, dap commands and sun
plan and the interdependence of these data products as well as the
operational constraints and safety measures we each have in place.
***********
SUSIM
***********
SUSIM performed its daily solar scans of the sun during periods of
predicted UARS telemetry.
***********
ACRIM
***********
ACRIM II continues to function normally and gather high quality data.
Re-processing is proceeding. Comparisons with ACRIM III demonstrate the
close agreement of both sets of observations relative to solar variable
features associated with solar maxima.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
HALOE continues successful scheduled operations. HALOE was powered
off beginning of day February 6, 2001. All engineering parameters were
within nominal operating region and science data products continue to be of
excellent quality. During this past operational period (26jan - 06feb) the
telescope field of view, for Sunset events viewing above 60 S, was set to
the center of the sun to reduce tracking uncertainties. This optimizes
measurements for studying Polar Mesospheric Clouds in the southern
hemisphere summer.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG participated in a meeting with the FOT off-line engineers
and Brian Boyle from SOLSTICE to discuss SOLSTICE operations.
Daily processing of instrument Daily Activity Plans continues
with no problems. Instrument operations schedule is now complete
to 15 March 2001.
***********
CDHF
***********
UARS CDHF REHOST:
- We plan to go operational on the REHOST on Wednesday February 14,2001.
Level 0, Quicklook and NRT will be transitioned over on that day.
If you use any of these products be sure you have an account on the REHOST
system which is set up to receive this data. If you do not please contact
PAC.
- Daily production processing will begin on this day also.
- PEM V3.32 (JOBS 1 and 2) have both tested successfully on the
Rehost System. Still waiting validation results from the PEM group.
STATUS:
-"Hold" on the HALOE 4.82 (jobs 2 - 4) due to a day gap
in the Auxiliary (SOL) data. The Solar group has been notified by the
MPG and also the Production group regarding the day gap. JOB 1 continues
to be processed against the current LEVEL0 files. The HALOE PI has been
notified also.
- All WINDII V11 (LEVEL3) data created from start of mission to 31-Aug-1997
(UARS DAYS 54 - 2169) that were requested by the WINDII group has been
loaded into the DAAC Transfer Tables. DAAC has been notified.
- Data created in the month of December, 2000 will be previewed and loaded
into the Daac Transfer Tables February, 2001.
***********
FOT
***********
REPORTING PERIOD
This report covers 27 January 2001 (Orbit 51271, GMT Day 027) through 02
February 2001 (Orbit 51375, GMT Day 033).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 028/1250 51293 Power On
The beta angle ranged between the angles of 54.0 degrees to 32.9 degrees for
this report period. The beta angle is now decreasing towards a minimum angle
of 02 degrees on 09 February 2001 (DOY 040).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.9 and 1.1 degrees C. Additional battery performance data for each
battery on 26 January 2001 and 02 February 2001 are provided in Appendix D.
The clock error ranged from ~13.3 to 11.2 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 02
¥ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 20
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 05
Ten (10) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/315 - 2287.496493
· Transponder B: 00/314 - 2287.498584
GROUND SYSTEM OPERATION
Five (5) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 01/032 show the observatory data loss to be 300 hours, 13
minutes, 27.048 seconds. This shows an overall increase of 8.272 seconds since
the last report period This is a 0.39 percent data loss, which equals a 99.61
percent data capture for the mission.
The increase of 8.272 seconds occurred over one (1) day:
· 8.272 seconds on Day 027 due to a Data Hit
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
On 27 January 2001 an ATC halt anomaly was experienced on UARS due to bad
safety load. The FOT caught the problem immediately and jumped the ATC pointer
to the proper location. Two SSPP commands were missed due to the halt, which
caused the SSPP to stop tracking for one orbit. Investigation found that the
safety load showed the proper file size on DFLX before load blocking, but load
was empty after blocking. System administrators and CMF personnel continue to
look into the load blocking anomaly and the DFLX system. A bad hard disk on
DFLX is the suspected cause of the anomaly and will be replaced next week.
After flying through the high beta period with the TAM as the only input to
UFLTR, UARS accrued larger residuals in the FSS measurements. On 31 January
2001 a new TAM alignment was uplinked to alleviate this problem, but still did
not seem to bring the FSS residuals down. The FSS was therefore put into
update mode to try and take out this attitude error. The FSS residuals came
down steadily as expected and UARS remains in the TAM and FSS configuration.
MEETINGS
The FOT attended an ESMO Anomaly Meeting on 30 January 2001.
The FOT attended a UARS Monthly Status Meeting on 31 January 2001.
Transitioned Honeywell employees attended the IPMD Training Meeting on 02
February 2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 09 February 2001 (DOY 040). This
maneuver will be a Reverse to Forward Flight maneuver (Southern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 02 February 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
51272 027/0232 01-057 OBC Clk Adj 20.6 to 0.00 msec/day
51278 027/1249 01-058 ATC ATC jump to location 95
51293 028/1250 01-061 HALOE Power ON
51298 028/1936 01-062 ATC NOOPS D-029
51299 028/2142 01-063 OBC Clk Adj 0.00 to 20.6 msec/day
51304 029/0551 01-062 NBTR NOOPS D-029
51310 029/1614 01-065 ACS TAM Alignment
51312 029/1755 01-062 NBTR NOOPS D-029
51312 029/1836 01-064 NBTR NOOPS D-030
51313 029/1939 01-066 ACS MAGMOD time update
51319 030/0552 01-064 NBTR NOOPS D-030
51324 030/1435 01-067 HALOE Adjust stow position
51324 030/1435 01-068 OBC Clk Adj 20.6 to 0.00 msec/day
51327 030/1836 01-069 NBTR NOOPS D-031
51334 031/0554 01-069 NBTR NOOPS D-031
51339 031/1436 01-070 ACS FSS in USE Mode
51342 031/1836 01-071 NBTR NOOPS D-032
51349 032/0639 01-071 NBTR NOOPS D-032
51357 032/1800 01-072 NBTR NOOPS D-033
51363 033/0415 01-072 NBTR NOOPS D-033
51364 033/0557 01-072 NBTR NBTR PBK/RTS-58
51372 033/1802 01-073 NBTR NOOPS D-034
___________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 02 February 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
HATBAT Y-HI HALOE Bolometer Area Temp 51315-51330
HASMIRRT Y-HI HALOE Secondary Mirror Temp 51322
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 02 February 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
51272 027/0236 Clk Adj 20.6 to 0.00 msec/day
51299 028/2145 Clk Adj 0.00 to 20.6 msec/day
51324 030/1440 Clk Adj 20.6 to 0.00 msec/day
___________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 02 February 2001
26 Jan 2001 - GMT Day 026 Beta = 58.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- ------- --------- ----- ------- -----
1 +212.8/+112.0 -3.05/-4.62 +0.0/+0.0 13.3 0.000 0.0
2 +11.2/-28.0 +4.19/+2.92 +24.0/-14.4 26.2 1.040 21.0
3 +16.8/-33.6 +3.23/+1.97 +23.2/-14.0 26.2 1.040 20.5
02 Feb 2001 - GMT Day 033 Beta = 32.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ---------- ------ ------- -----
1 +39.2/+0.0 -3.68/-4.62 +0.0/+0.0 13.0 0.000 0.0
2 +11.2/-22.4 +5.47/+3.87 +28.8/-14.4 25.9 1.040 20.3
3 +22.4/-22.4 +4.51/+2.92 +28.4/-15.2 25.9 1.030 19.9
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
---- ---- ---------
-no changes V/T 5
___________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 02 February 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-005 FOT 51276 027/0917 N/A Load u027z101 cont. no cmds.
01-006 FOT 51280 027/1514 N/A SSPP missed one star track
01-007 FOT 51310 029/1614 N/A CMS fails to respond to ftr
01-008 FOT 51334 031/0554 23548 TPF not receiving data
01-009 FOT 51362 033/0236 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-053 FOT 50168 319/1113 23157 Late Acquisition
00-054 FOT 50280 326/2147 23187 RESET PTP /2:39 32K
00-056 FOT 50584 347/0502 23302 Premature LOS
01-002 FOT see list Jan 01 14170 Generic Late Acquisition
01-003 FOT 51111 016/0856 N/A AP Crashed/disk crash
01-004 FOT 51220 023/1606 23512 Late Acquisition/4:02 non-rec
01-008 FOT 51334 031/0554 23548 TPF not receiving data
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-005 FOT 51276 027/0917 N/A Load u027z101 cont. no cmds.
01-006 FOT 51280 027/1514 N/A SSPP missed one star track
01-007 FOT 51310 029/1614 N/A CMS fails to respond to ftr
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-005 FOT 51276 027/0917 N/A Load u027z101 cont. no cmds.
01-006 FOT 51280 027/1514 N/A SSPP missed one star track
01-007 FOT 51310 029/1614 N/A CMS fails to respond to ftr
01-009 FOT see list Feb 01 14170 Generic Late Acquisition
SEVNTFW01/05